Life insurance salesman meets his prospective customers separately. At each meeting, he makes the same effort to persuade the prospective client to be insured. The vendor's understanding after a long term in the insurance business is that the probability of persuading the customer to be insured (probability of success) is 0,1. What is the probability in 4 trials to insure one of them?
1
Expert's answer
2020-09-29T15:56:36-0400

Given that,


Probability of success, p = 0.1

Probability of failure, q = 1 - p = 1 - 0.1 = 0.9

Number of total outcomes n = 4

Number of favorable outcomes x = 1


By using the formula for binomial distribution,


P = "nCx\ *\ p^x * q^{n-x}"


P = "4C1\ *\ 0.1^1\ *\ 0.9^{4-1}"


P = 0.2916


Hence, the probability that in 4 trials of meeting 1 client gets insured is 0.2916